Factor and simplify cos^2x-sin^2x cos^2x
Answer by lwsshak3(11628)

Factor and simplify
cos^2x-sin^2x cos^2x
=cos^2x(1-sin^2x)
=cos^2x(cos^2x)
=cos^4x